A parallel beam of light falls normally on one phase of a small angled glass prism. Part of the beam

is refracted at the second phase and is deviated through an angle of 1.8o

. Another part is reflected

at the second phase and emerges at the first making an angle of 9.2o

with the normal. Calculate:

(i) The refractive index of the glass.

(ii) The angle of the prism.


1
Expert's answer
2021-04-01T18:41:41-0400

Answer

We know according to optics when A parallel beam of light falls normally on one phase of a small angled glass.

So we can write

"\Delta=\theta(n-1)" ----(1)

where "\Delta" is the deviation angle, θ is angle of the prism, n is the refractive index of the glass of given data.

According to condition, we have

"i_1=0" °

"r_1-i_1=1.8\u00b0"

"i_2-r_2=9.2\u00b0"

So we have

r2=r1=1.8°


The deviation angle is equal to

"\u03b4=9.2+1.8=11\u00b0"


According to Snellus's law, we get


"\frac {\sin {i_2}}{\sin {r_2}}=n----- (2)"

Using (1) we get


n=1.96


Using (1), we get as

11.46



(So The refractive index of the glass is equal to 1.96.

And also he angle of the prism is equal to 11.46